1.
the length of the sides of a triangle are x+2,2x+3 and 5x. If perimeter is 21 find the lengths of the sides?
the length of the shorts side is?
the length of the longest side is?
the length of the intermediate is?

Answer by algebrahouse.com(1659)

x + 2 + 2x + 3 + 5x = 21 {perimeter of triangle is 21}
8x + 5 = 21 {combined like terms}
8x = 16 {subtracted 5 from each side}
x = 2 {divided each side by 8}

x + 2 = 4 {substituted 2, in for x, into x + 2}
2x + 3 = 7 {substituted 2, in for x, into 2x + 3}
5x = 10 {substituted 2, in for x, into 5x}

short side is 4
longest side is 10
intermediate side is 7

For more help from me, visit: www.algebrahouse.com